OpenStudy (anonymous):

find the greatest common factor of 1/3x^2+1/3x-4

OpenStudy (zehanz):

\[\frac{ 1 }{ 3 } x^2+\frac{ 1 }{ 3 } x-4 \]You can factor out 1/3:\[\frac{ 1 }{ 3 }(x^2+x-12)\]1/3 is actually a factor of 4, because 1/3 * 12 = 4